Fred leaves his house for work, which is exactly 2 miles away. After Fred has traveled 2/3 mile he must turn around to pick up some papers. After he travels 1/6 mile back home, how far away from work is he?
Answer by ankor@dixie-net.com(22740)

Fred leaves his house for work, which is exactly 2 miles away. After Fred has traveled 2/3 mile he must turn around to pick up some papers. After he travels 1/6 mile back home, how far away from work is he?
:
2 - (2/3 - 1/6)
Use 6ths
12/6 - (4/6 - 1/6) =
12/6 - 3/6 = 9/6 = 1.5 mi from work